(Calculating financial ratios) The balance sheet and income statement for the J·P. Robard Mig. Company are as follows:囲. Calculate the following ratios: Current ratio Times interest earned Inventory turnover Total asset turnover Operating profit margin Operating return on assets Debt ratio Average collection period Fixed asset turnover Return on equityP. Robard Mfg., Inc. J. Balance Sheet ($000) $500 2,000 1,000 $3,500 4,500 $8,000 Cash Accounts receivable Inventories Current assets Net fixed assets Total assets Accounts payable Accrued expenses Short-term notes payable $1,100 600 300 $2,000 2,000 4,000 $8,000 Current liabilities Long-term debt Owners equity Total liabilities and owners equityJ. P. Robard Mfg., Ind Income Statement ($000) $8,000 (3,300) $4,700 (3,000) $1,700 (367) $1,333 (533) $800 Net sales (all credit) Cost of goods sold Gross profit Net operating income Earnings before taxes Interest expense Income taxes (40%) Net income
Answer #1

Current Ratio = Current Assets/Current liabilities = 3500/2000 = 1.75
Times interest Earned =(Net operating Income) EBIT/Interest = 1700/367 = 4.63
Inventory Turnover = C0GS/Inventory = 3300/1000 = 3.3
Total Assets Turnover = Sales/Total Assets = 8000/8000 = 1
Operating Profit Margin = Operating income/Sales = 1700/8000 = 21.25% or 0.2125
Operating Return on Assets = Net Operating Income/Total Assets = 1700/8000 = 21.25% or 0.2125
Debt Ratio =Total Debt/Total Assets = 4000/8000 = 0.5
Average Collection Period = 365 * Account Receivables /COGS = 365* 2000/3300 = 221.21
Fixed Assets Turnover = Sales/Total Fixed Assets = 8000/4500 = 1.78
Return on Equity = Net Income/Equity = 800/4000 = 0.2

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well